OpenStudy (anonymous):

Which is the solution set to the given inequality? |x 4| < 12

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

rewrite the question.

OpenStudy (chosenmatt):

divide both sides by 4 i think

OpenStudy (haseeb96):

what is sign between x and 4

OpenStudy (anonymous):

@Dylan0318

OpenStudy (anonymous):

-

OpenStudy (anonymous):

i agree with @chosenmatt

OpenStudy (chosenmatt):

yay someone agrees with me!

OpenStudy (anonymous):

then -12<x-4<12 implies -8<x<16 is the required solution

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i.e. (-8,16) this interval is the solution

OpenStudy (anonymous):

Thanks

OpenStudy (haseeb96):

there will be two solution for this x-4 =12 and x-4=-12 x=12+4 and x=-12+4 x=16 and x=-8

OpenStudy (haseeb96):

this is the correct answer

OpenStudy (anonymous):

the sign is less then, not equal. @Haseeb96 abd @chosenmatt

OpenStudy (chosenmatt):

oh

OpenStudy (anonymous):

and it is x-4 not x*4

OpenStudy (chosenmatt):

k

OpenStudy (haseeb96):

i know we can put it now this was my way of teaching now x<16 and x >-8

OpenStudy (anonymous):

I am giving you a general formula |a-b|< c implies -c<a-b<c